the perimeter of a rectangle is twice the sum of its length and its width. the perimeter is 40 meters and its length is 2 meters more than twice its width. what is the length?
Answer by Deina(147)

P = 40
P = 2(L + W)
L = 2W + 2
P = 2(2W + 2 + W)
P = 2(3W + 2)
P = 6W + 4
40 = 6W + 4 | subtract 4 from each side:
36 = 6W | divide each side by 6
W = 6

L = 2*6 + 2 = 14
2(14 + 6) = 40 and bob's your uncle!